# Gradient Estimation with Discrete Stein Operators

### v7

@@ 5 Related Work
As we have seen in Section 2, there is a long history of designing CVs for REINFORCE estimators using “baselines” (Weaver and Tao, 2001; Bengio et al., 2013; Ranganath et al., 2014; Mnih and Gregor, 2014). Recent progress is mostly driven by leave-one-out (Salimans and Knowles, 2014; Mnih and Rezende, 2016; Kool et al., 2019; Richter et al., 2020) and sample-dependent baselines (Paisley et al., 2012; Gu et al., 2016; Titsias and Shi, 2022; Tucker et al., 2017; Grathwohl et al., 2018). REBAR (Tucker et al., 2017) constructs the baseline through continuous relaxation of the discrete distribution (Jang et al., 2017; Maddison et al., 2017) and applies reparameterization gradients to the correction term. As a result REBAR uses three evaluations of $f$ for each $x^{(k)}$ instead of the usual single evaluation (see Appendix A for a detailed explanation). The RELAX (Grathwohl et al., 2018) estimator generalizes REBAR by noticing that their continuous relaxation can be replaced with a free-form CV. However, in order to get strong performance, RELAX still includes the continuous relaxation in their CV and only adds a small deviation to it. Therefore, RELAX also uses three evaluations of $f$ for each $x^{(k)}$ and is usually considered more expensive than other estimators.
@@
An attractive property of the RODEO estimator is that it incorporates information from neighboring states thanks to the Stein operator while avoiding additional costly $f$ evaluations thanks to the learned surrogate functions $h_{k}$. Estimators based on analytic local expectation (Titsias and Lázaro-Gredilla, 2015; Tokui and Sato, 2017) and GO gradients (Cong et al., 2019) also use neighborhood information but only at the cost of many additional target function evaluations. In fact, the local expectation gradient (Titsias and Lázaro-Gredilla, 2015) can be viewed as a Stein CV adjustment RODEO with a Gibbs Stein operator and the target function $f$ used directly instead of the surrogate $h$. The downside of these approaches is that $f$ must be evaluated $Kd$ times per training step instead of $K$ times as in RODEO, a prohibitive cost when $f$ is expensive and $d\geq 200$ as in Section 6.
@@
Prior work has also studied variance reduction methods based on sampling without replacement (Kool et al., 2020) and antithetic sampling (Yin and Zhou, 2019; Dong et al., 2020; Dimitriev and Zhou, 2021; Dong et al., 2021) for gradient estimation. DisARM (Dong et al., 2020; Yin et al., 2020) was shown to outperform RLOO estimators when $K=2$ and ARMS (Dimitriev and Zhou, 2021) generalizes DisARM to $K>2$.
@@
Stein operators for continuous distributions have also been leveraged for effective variance reduction in a variety of learning tasks (Assaraf and Caffarel, 1999; Mira et al., 2013; Oates et al., 2017; Barp et al., 2018; South et al., 2018; Si et al., 2020) including gradient estimation (Liu et al., 2017; Jankowiak and Obermeyer, 2018). In particular, the gradient estimator of Liu et al. (2017) is based on the Langevin Stein operator (Gorham and Mackey, 2015) for continuous distributions and coincides with the continuous counterpart of RELAX (Grathwohl et al., 2018). In contrast, our approach considers discrete Stein operators for Monte Carlo estimation in discrete distributions with exponentially large state spaces. Recently, Parmas and Sugiyama (2021, App. E.4) used a probability flow perspective to characterize all unbiased gradient estimators satisfying a mild technical condition; our estimators fall into this broad class but were not specifically investigated.
@@ 6.1 Training Bernoulli VAEs
Following Dong et al. (2020) and Titsias and Shi (2022), we conduct experiments on training variational auto-encoders (Kingma and Welling, 2014; Rezende et al., 2014) (VAEs) with Bernoulli latent variables. VAEs are models with a joint density $p_{\theta}(y,x)$, where $x$ is the latent variable and $\theta$ denotes model parameters. They are typically learned through maximizing the evidence lower bound (ELBO) $\mathbb{E}_{q_{\eta}(x|y)}[f(x)]$ for an auxiliary inference network $q_{\eta}(x|y)$ and $f(x)\triangleq\log p_{\theta}(y,x)-\log q_{\eta}(x|y)$. In our experiments, $p(x)$ and $q_{\eta}(x|y)$ are high-dimensional distributions where each dimension of the random variable is an independent Bernoulli. Since exact gradient computations are intractable, we will use gradient estimators to learn the parameters $\eta$ of the inference network. The VAE architecture and training experimental setup follows Titsias and Shi (2022), and details are given in Appendix D. The dimensionality of the latent variable $x$ is $d=200$. The functions $H$ 13 and $H^{*}$ 14 share a neural network architecture with two output units and a single hidden layer with 100 units. For the numerical stability and variance reasons discussed in Section 3, we use the Gibbs Stein operator 4 as a default choice in our experiments, but we revisit this choice in Section 6.3.
We consider the MNIST (LeCun et al., 1998), Fashion-MNIST (Xiao et al., 2017) and Omniglot (Lake et al., 2015) datasets using their standard train, validation, and test splits. We use both binary and continuous VAE outputs ($y$) as in Titsias and Shi (2022). In the binary output setting, data are dynamically binarized, and the Bernoulli likelihood is used; in the continuous output setting, data are centered between $[-1,1]$, and the Gaussian likelihood with learnable diagonal covariance parameters is used.
@@
$\boldsymbol{K=2}$   In the first set of experiments we focus on the most common setting of $K=2$ sample points and compare our variance reduction method to its counterparts including DisARM (Dong et al., 2020) and Double CV (Titsias and Shi, 2022). Results for RLOO are omitted since it is consistently outperformed by Double CV (see Titsias and Shi, 2022). Table 2 shows, on all three datasets, RODEO achieves the best training ELBOs. In Figure 1 (left), we plot the gradient variance and average training ELBOs against training steps for all estimators on dynamically binarized MNIST. RODEO outperforms DisARM and Double CV by a large margin in gradient variance.
@@
$\boldsymbol{K=3}$   In the second set of experiments we compare RELAX (Grathwohl et al., 2018), which uses three evaluations of $f$ per training step, with RODEO, Double CV, and ARMS (Dimitriev and Zhou, 2021) for $K=3$. Figure 1 (right) and Table 2 demonstrate that RODEO outperforms the three previous methods and generally leads to the lowest variance. Although RELAX was often observed to have very strong performance in prior work (Dong et al., 2020; Titsias and Shi, 2022), our results in Figure 1 suggest that, for dynamically binarized datasets, much larger gains can be achieved by using the same number of function evaluations in other estimators.
@@ 6.2 Training hierarchical Bernoulli VAEs
To investigate the performance of RODEO when scaling to hierarchical discrete latent variable models, we follow DisARM (Dong et al., 2020; Yin and Zhou, 2019) to train VAEs with 2/3/4 stochastic layers, each of which consists of $200$ Bernoulli variables. We set $K=2$ and compare our estimator with DisARM and Double CV on dynamically binarized MNIST, Fashion-MNIST, and Omniglot. For each stochastic layer, we use a different CV network which has the same architecture as those in our VAE experiments from the previous section. More details are presented in Appendix D.
@@ 6.3 Ablation study
Impact of surrogate functions   To tease apart the benefits of our new surrogate functions 13 and the remaining RODEO components, we replace the surrogate function $c_{\phi}(z)$ in RELAX (Grathwohl et al., 2018) with our surrogates, which only requires a single evaluation of $f$ per $x^{(k)}$ (see Appendix A for more details). Figure 3(c) compares the performance of RODEO and this modified RELAX on binarized MNIST. Since the surrogate functions are matched, the consistent improvements over modified RELAX can be attributed to the Stein and double CV components of RODEO. In Section C.2, we also experiment with increasing the complexity of $H$ and $H^{*}$ by using two hidden layers instead of a single hidden layer in the MLP. We did not observe significant improvements in variance reduction.
@@ 7 Conclusions, Limitations, and Future Work
One potential drawback of our surrogate function constructions 13 and 14 is the need to evaluate an auxiliary function ($H$ or $H^{*}$) at $K-1$ locations. This cost can be comfortably borne when $H$ and $H^{*}$ are much cheaper to evaluate than $f$, such as in the ResNet VAE example in Section C.1 and many large VAE models used in practice (Vahdat and Kautz, 2020). And, in our experiments with the most common sample sizes, the runtime of RODEO was no worse than that of RELAX. To obtain a more favorable cost for large $K$, one could employ alternative surrogates that require only a constant number of auxiliary function evaluations, e.g.,
@@ Appendix A Sample-dependent Baselines in REBAR and RELAX
REBAR (Tucker et al., 2017) introduces indirect independence on $x^{(k)}$ in $b_{k}$ through the continuous reparameterization $x=H(z),z^{(k)}\sim q_{\eta}(z|x=x^{(k)})$, where $z$ is a continuous variable and $H$ is an argmax-like thresholding function. Specifically, $b_{k}=f(\sigma_{\lambda}(z^{(k)}))$, where $\sigma_{\lambda}$ is a continuous relaxation of $H$ controlled by the parameter $\lambda$. The correction term decomposes into two parts:
@@
Both parts can be estimated with the reparameterization trick (Kingma and Welling, 2014; Rezende et al., 2014; Titsias and Lázaro-Gredilla, 2014) which often has low variance. The RELAX (Grathwohl et al., 2018) estimator generalizes REBAR by noticing that $f(\sigma_{\lambda}(z))$ can be replaced with a free-form differentiable function $c_{\phi}(z)$. However, RELAX still relies on parameterizing $c_{\phi}(z)$ as $f(\sigma_{\lambda}(z))+r_{\theta}(z)$ to achieve strong performance, as noted in Dong et al. (2020).
@@ C.1 Wall clock time comparison with RLOO
Besides necessary target function evaluations, the RODEO estimator comes with the additional cost of evaluating the neural network-based $H,H^{*}$. Therefore, RODEO is most suited to the problems where the cost of evaluating $f$ dominates that of evaluating $H,H^{*}$. This is often the case in practice. For example, state-of-the-art variational autoencoders (e.g., Vahdat and Kautz, 2020) are often built on expensive neural architectures such as deep residual networks (ResNets). Here, to demonstrate the practical advantage of our method as the complexity of $f$ grows, we replace the two-layer MLP VAEs used in previous experiments with a ResNet VAE (architecture shown in Table 3), while the neural network used by $H,H^{*}$ remains a single-layer MLP with 100 hidden units. We then compare the wall clock performance of RODEO with RLOO. The latent variables in this experiment remain binary and have 200 dimensions.
@@ Appendix D Experimental Details
Our implementation is based on the open-source code of DisARM (Dong et al., 2020) (Apache license) and Double CV (Titsias and Shi, 2022) (MIT license). Our figures display 1M training steps and our tables report performance after 1M training steps to replicate the experimental settings of DisARM (Dong et al., 2020).
@@ D.1 Details of VAE experiments
VAEs are models with a joint density $p(y,x)=p(y|x)p(x)$, where $x$ denotes the latent variable. $x$ is assigned a uniform factorized Bernoulli prior. The likelihood $p_{\theta}(y|x)$ is parameterized by the output of a neural network with $x$ as input and parameters $\theta$. The VAE has two hidden layers with $200$ units activated by LeakyReLU with the coefficient $0.3$. To optimize the VAE we use Adam with base learning rate $10^{-4}$ for non-binarized data and $10^{-3}$ for dynamically binarized data, except for binarized Fashion-MNIST we decreased the learning rate to $3\times 10^{-4}$ because otherwise the training is very unstable for all estimators. We use Adam with the same learning rate $10^{-3}$ for adapting our control variate network in all experiments. The batch size is $100$. The settings of other estimators are kept the same with Titsias and Shi (2022).
@@ D.2 Details of hierarchical VAE experiments
We optimize the hierarchical VAE using Adam with base learning rate $10^{-4}$. Our control variate network is optimized using Adam with learning rate $10^{-3}$. Settings of training multilayer VAEs are kept the same with Dong et al. (2020), except that we do not optimize the prior distribution of the VAE hidden layer and use a larger batch size $100$.
@@ Appendix E Additional Results
In this section, we measure test set performance using 100 test points and the marginal log-likelihood bound of Burda et al. (2015), which provides a tighter estimate of marginal log likelihood than the ELBO. Throughout, we call this the “test log-likelihood bound.”

### v8

@@ 5 Related Work
As we have seen in Section 2, there is a long history of designing CVs for REINFORCE estimators using “baselines” [64, 7, 45, 37]. Recent progress is mostly driven by leave-one-out [49, 38, 30, 48] and sample-dependent baselines [43, 22, 60, 62, 20]. REBAR [62] constructs the baseline through continuous relaxation of the discrete distribution [26, 35] and applies reparameterization gradients to the correction term. As a result REBAR uses three evaluations of $f$ for each $x^{(k)}$ instead of the usual single evaluation (see Appendix A for a detailed explanation). The RELAX [20] estimator generalizes REBAR by noticing that their continuous relaxation can be replaced with a free-form CV. However, in order to get strong performance, RELAX still includes the continuous relaxation in their CV and only adds a small deviation to it. Therefore, RELAX also uses three evaluations of $f$ for each $x^{(k)}$ and is usually considered more expensive than other estimators.
@@
An attractive property of the RODEO estimator is that it incorporates information from neighboring states thanks to the Stein operator while avoiding additional costly $f$ evaluations thanks to the learned surrogate functions $h_{k}$. Estimators based on analytic local expectation [59, 61] and GO gradients [11] also use neighborhood information but only at the cost of many additional target function evaluations. In fact, the local expectation gradient [59] can be viewed as a Stein CV adjustment RODEO with a Gibbs Stein operator and the target function $f$ used directly instead of the surrogate $h$. The downside of these approaches is that $f$ must be evaluated $Kd$ times per training step instead of $K$ times as in RODEO, a prohibitive cost when $f$ is expensive and $d\geq 200$ as in Section 6.
@@
Prior work has also studied variance reduction methods based on sampling without replacement [31] and antithetic sampling [68, 14, 13, 15] for gradient estimation. DisARM [14, 69] was shown to outperform RLOO estimators when $K=2$ and ARMS [13] generalizes DisARM to $K>2$.
@@
Stein operators for continuous distributions have also been leveraged for effective variance reduction in a variety of learning tasks [2, 36, 41, 5, 54, 52] including gradient estimation [34, 27]. In particular, the gradient estimator of Liu et al. [34] is based on the Langevin Stein operator [19] for continuous distributions and coincides with the continuous counterpart of RELAX [20]. In contrast, our approach considers discrete Stein operators for Monte Carlo estimation in discrete distributions with exponentially large state spaces. Recently, Parmas and Sugiyama [44, App. E.4] used a probability flow perspective to characterize all unbiased gradient estimators satisfying a mild technical condition; our estimators fall into this broad class but were not specifically investigated.
@@ 6.1 Training Bernoulli VAEs
Following Dong et al. [14] and Titsias and Shi [60], we conduct experiments on training variational auto-encoders [29, 47] (VAEs) with Bernoulli latent variables. VAEs are models with a joint density $p_{\theta}(y,x)$, where $x$ is the latent variable and $\theta$ denotes model parameters. They are typically learned through maximizing the evidence lower bound (ELBO) $\mathbb{E}_{q_{\eta}(x|y)}[f(x)]$ for an auxiliary inference network $q_{\eta}(x|y)$ and $f(x)\triangleq\log p_{\theta}(y,x)-\log q_{\eta}(x|y)$. In our experiments, $p(x)$ and $q_{\eta}(x|y)$ are high-dimensional distributions where each dimension of the random variable is an independent Bernoulli. Since exact gradient computations are intractable, we will use gradient estimators to learn the parameters $\eta$ of the inference network. The VAE architecture and training experimental setup follows Titsias and Shi [60], and details are given in Appendix D. The dimensionality of the latent variable $x$ is $d=200$. The functions $H$ 13 and $H^{*}$ 14 share a neural network architecture with two output units and a single hidden layer with 100 units. For the numerical stability and variance reasons discussed in Section 3, we use the Gibbs Stein operator 4 as a default choice in our experiments, but we revisit this choice in Section 6.3.
We consider the MNIST [33], Fashion-MNIST [66] and Omniglot [32] datasets using their standard train, validation, and test splits. We use both binary and continuous VAE outputs ($y$) as in Titsias and Shi [60]. In the binary output setting, data are dynamically binarized, and the Bernoulli likelihood is used; in the continuous output setting, data are centered between $[-1,1]$, and the Gaussian likelihood with learnable diagonal covariance parameters is used.
@@
$\boldsymbol{K=2}$   In the first set of experiments we focus on the most common setting of $K=2$ sample points and compare our variance reduction method to its counterparts including DisARM [14] and Double CV [60]. Results for RLOO are omitted since it is consistently outperformed by Double CV [see 60]. Table 2 shows, on all three datasets, RODEO achieves the best training ELBOs. In Figure 1 (left), we plot the gradient variance and average training ELBOs against training steps for all estimators on dynamically binarized MNIST. RODEO outperforms DisARM and Double CV by a large margin in gradient variance.
@@
$\boldsymbol{K=3}$   In the second set of experiments we compare RELAX [20], which uses three evaluations of $f$ per training step, with RODEO, Double CV, and ARMS [13] for $K=3$. Figure 1 (right) and Table 2 demonstrate that RODEO outperforms the three previous methods and generally leads to the lowest variance. Although RELAX was often observed to have very strong performance in prior work [14, 60], our results in Figure 1 suggest that, for dynamically binarized datasets, much larger gains can be achieved by using the same number of function evaluations in other estimators.
@@ 6.2 Training hierarchical Bernoulli VAEs
To investigate the performance of RODEO when scaling to hierarchical discrete latent variable models, we follow DisARM [14, 68] to train VAEs with 2/3/4 stochastic layers, each of which consists of $200$ Bernoulli variables. We set $K=2$ and compare our estimator with DisARM and Double CV on dynamically binarized MNIST, Fashion-MNIST, and Omniglot. For each stochastic layer, we use a different CV network which has the same architecture as those in our VAE experiments from the previous section. More details are presented in Appendix D.
@@ 6.3 Ablation study
Impact of surrogate functions   To tease apart the benefits of our new surrogate functions 13 and the remaining RODEO components, we replace the surrogate function $c_{\phi}(z)$ in RELAX [20] with our surrogates, which only requires a single evaluation of $f$ per $x^{(k)}$ (see Appendix A for more details). Figure 4(c) compares the performance of RODEO and this modified RELAX on binarized MNIST. Since the surrogate functions are matched, the consistent improvements over modified RELAX can be attributed to the Stein and double CV components of RODEO. In Section C.2, we also experiment with increasing the complexity of $H$ and $H^{*}$ by using two hidden layers instead of a single hidden layer in the MLP. We did not observe significant improvements in variance reduction.
@@ 7 Conclusions, Limitations, and Future Work
One potential drawback of our surrogate function constructions 13 and 14 is the need to evaluate an auxiliary function ($H$ or $H^{*}$) at $K-1$ locations. This cost can be comfortably borne when $H$ and $H^{*}$ are much cheaper to evaluate than $f$, such as in the ResNet VAE example in Section C.1 and many large VAE models used in practice [63]. And, in our experiments with the most common sample sizes, the runtime of RODEO was no worse than that of RELAX. To obtain a more favorable cost for large $K$, one could employ alternative surrogates that require only a constant number of auxiliary function evaluations, e.g.,
@@ Appendix A Sample-dependent Baselines in REBAR and RELAX
REBAR [62] introduces indirect independence on $x^{(k)}$ in $b_{k}$ through the continuous reparameterization $x=H(z),z^{(k)}\sim q_{\eta}(z|x=x^{(k)})$, where $z$ is a continuous variable and $H$ is an argmax-like thresholding function. Specifically, $b_{k}=f(\sigma_{\lambda}(z^{(k)}))$, where $\sigma_{\lambda}$ is a continuous relaxation of $H$ controlled by the parameter $\lambda$. The correction term decomposes into two parts:
@@
Both parts can be estimated with the reparameterization trick [29, 47, 58] which often has low variance. The RELAX [20] estimator generalizes REBAR by noticing that $f(\sigma_{\lambda}(z))$ can be replaced with a free-form differentiable function $c_{\phi}(z)$. However, RELAX still relies on parameterizing $c_{\phi}(z)$ as $f(\sigma_{\lambda}(z))+r_{\theta}(z)$ to achieve strong performance, as noted in Dong et al. [14].
@@ C.1 Wall clock time comparison with RLOO
Besides necessary target function evaluations, the RODEO estimator comes with the additional cost of evaluating the neural network-based $H,H^{*}$. Therefore, RODEO is most suited to the problems where the cost of evaluating $f$ dominates that of evaluating $H,H^{*}$. This is often the case in practice. For example, state-of-the-art variational autoencoders [e.g., 63] are often built on expensive neural architectures such as deep residual networks (ResNets). Here, to demonstrate the practical advantage of our method as the complexity of $f$ grows, we replace the two-layer MLP VAEs used in previous experiments with a ResNet VAE (architecture shown in Table 3), while the neural network used by $H,H^{*}$ remains a single-layer MLP with 100 hidden units. We then compare the wall clock performance of RODEO with RLOO. The latent variables in this experiment remain binary and have 200 dimensions.
@@ Appendix D Experimental Details
Our implementation is based on the open-source code of DisARM [14] (Apache license) and Double CV [60] (MIT license). Our figures display 1M training steps and our tables report performance after 1M training steps to replicate the experimental settings of DisARM [14].
@@ D.1 Details of VAE experiments
VAEs are models with a joint density $p(y,x)=p(y|x)p(x)$, where $x$ denotes the latent variable. $x$ is assigned a uniform factorized Bernoulli prior. The likelihood $p_{\theta}(y|x)$ is parameterized by the output of a neural network with $x$ as input and parameters $\theta$. The VAE has two hidden layers with $200$ units activated by LeakyReLU with the coefficient $0.3$. To optimize the VAE we use Adam with base learning rate $10^{-4}$ for non-binarized data and $10^{-3}$ for dynamically binarized data, except for binarized Fashion-MNIST we decreased the learning rate to $3\times 10^{-4}$ because otherwise the training is very unstable for all estimators. We use Adam with the same learning rate $10^{-3}$ for adapting our control variate network in all experiments. The batch size is $100$. The settings of other estimators are kept the same with Titsias and Shi [60].
@@ D.2 Details of hierarchical VAE experiments
We optimize the hierarchical VAE using Adam with base learning rate $10^{-4}$. Our control variate network is optimized using Adam with learning rate $10^{-3}$. Settings of training multilayer VAEs are kept the same with Dong et al. [14], except that we do not optimize the prior distribution of the VAE hidden layer and use a larger batch size $100$.
@@ Appendix E Additional Results
In this section, we measure test set performance using 100 test points and the marginal log-likelihood bound of Burda et al. [10], which provides a tighter estimate of marginal log likelihood than the ELBO. Throughout, we call this the “test log-likelihood bound.”